# Unified Framework for Geodesic Dynamics with Conservative, Dissipative, and GUP Effects

### v1

@@ Abstract
We derive generalized geodesic equations in curved spacetime that include conservative forces, dissipative effects, and quantum-gravity-motivated minimal-length corrections. Conservative interactions are incorporated through external vector potentials, while dissipative dynamics arise from an exponential rescaling of the particle Lagrangian. Quantum-gravity effects are introduced via Generalized Uncertainty Principle (GUP) deformed Poisson brackets in the Hamiltonian framework. We show that free-particle geodesics remain unaffected at leading order, but external potentials induce velocity-dependent corrections, implying possible violations of the equivalence principle. As an application, we analyze modified trajectories in Friedmann-Lemaitre-Robertson-Walker (FLRW) universes dominated by dust, radiation, stiff matter, and dark energy. Our results establish a unified approach to conservative, dissipative, and GUP-corrected geodesics, providing a framework to probe the interplay between external forces, spacetime curvature, and Planck-scale physics.
@@
General Relativity (GR) is built on two key ideas. The first is the equivalence principle, which says that gravity acts the same on all objects, no matter their shape or composition. The second is Mach’s principle, which states that matter shapes the geometry of spacetime, and this curvature then guides matter to move along the straightest possible paths, called geodesics and they are crucial for understanding phenomena such as cosmic horizons, dark energy, and gravitational lensing, as well as for studying observational effects like cosmic bubble collisions and theoretical frameworks such as the AdS/CFT correspondence [5, 6, 7, 8].
@@ III.1.2 Example for dissipative Lagrangian
From Eq.(8), we can obtain the dissipation exponent for the FLRW spacetime in the comoving frame with four-velocity $u^{\kappa}=(1,0,0,0)$. In this case,
@@
where $H(t)$ is the Hubble parameter. On choosing the dissipation function $Q(x^{\alpha})$ to be constant, the dissipation exponent becomes a function of time only, i.e. $\Gamma=\Gamma(t)$. In this case, the geodesic equation simplifies to
@@
where the term $\partial_{0}\Gamma=2H(t)Q(t)$. The physical origin of dissipation in this framework is encoded in the scalar function $Q(x^{\alpha})$. It acts as a phenomenological parameter that characterizes the exchange of energy between the particle and an external environment, such as a thermal bath, background radiation, or effective degrees of freedom that have been coarse-grained out. The coupling with the spacetime expansion, expressed through $\nabla_{\kappa}u^{\kappa}$, ensures that the dissipative effect depends on the local geometry and flow. The geodesic path will be affected by the values of the dissipation function $Q$, which is chosen to be constant here. For positive value ($Q>0$), it means the energy is dissipating in the surrounding, while negative values of $Q$ correspond to effective energy injection as in dissipative quintessence and inflationary scenarios [13]. The qualitative behavior of these set of differential equations, Eq.(31,32), through geodesic curve is represented in Fig.2. We show some plot of geodesic curves for different values of $Q$.
@@ IV The Geodesic equations under minimal length effect (GUP)
The existence of a minimal measurable length in quantum gravity requires a modification of the standard Heisenberg uncertainty principle. This leads to the Generalized Uncertainty Principle (GUP), which can be expressed as
@@
where $x$ and $p$ denote the position and its canonically conjugate momentum, respectively. The deformation parameter $\beta$ encodes the strength of quantum-gravity that became significant where the quantum gravitational effects are visible. The parameter $\beta$ typically arises in various approaches to quantum gravity, such as string theory, loop quantum gravity, and doubly special relativity. The dimension of $\beta$ is inverse of the square of Planck momentum as represented as $\beta=\beta_{0}/(m_{P}c)^{2}$, where $m_{P}$ is the Planck mass and $\beta_{0}$ is a dimensionless parameter. The value of $\beta_{0}$ is model-dependent and can be fixed or constrained through phenomenological studies in both quantum-mechanical and classical regimes.
@@
It should be emphasized that this procedure, taking the GUP-deformed commutator and promoting it directly to a Poisson bracket, has limitations. In particular, as discussed in [66], when taking the formal limit $\hbar\rightarrow 0$, the deformation term often diverges as $(1/\hbar)$. Strictly speaking, this signals an inconsistency in deriving the classical limit directly from the quantum commutator. On the other hand, it has been argued in [67] that although the factor $\hbar^{-1}$ diverges in the classical limit, the GUP parameter $\beta$ encodes the strength of quantum-gravity effects, where the dimensionless parameter $\beta_{0}$ can be chosen to be vanishingly small. This ensures that the overall factor $\beta_{0}\hbar^{-1}$ remains small but finite, thereby providing a consistent classical limit.
@@ IV.2 GUP-Corrected Geodesic Equation in the Presence of a Potential
We extended our analysis by incorporating minimal-length effects through a GUP-deformed symplectic structure at the classical level. Our results show that free-particle geodesics remain unaffected to leading order in the deformation parameter $\beta$, thereby preserving the equivalence principle in the absence of external interactions. However, when external potentials are present, the geodesic equation acquires velocity-dependent corrections proportional to $\beta$. These velocity-dependent modifications in the GUP-modified geodesic equation signal violations of the equivalence principle.
Together, these findings demonstrate that while pure geodesic motion is robust against Planck-scale modifications, the interplay between external forces, dissipation, and GUP corrections leads to a rich phenomenology, particularly in cosmological backgrounds. Our qualitative analysis of trajectories in dust, radiation, stiff matter, and dark-energy dominated universes illustrates how such effects can alter the geodesic paths at both the classical and quantum-gravity levels, as depicted in Figs. (1, 2, 3).

### v2

@@ Abstract
We derive generalized geodesic equations in curved spacetime that include conservative forces, dissipative effects, and quantum-gravity-motivated minimal-length corrections. Conservative interactions are incorporated through external vector potentials, while dissipative dynamics arise from an exponential rescaling of the particle Lagrangian. Phenomenological study of Quantum-gravity effects is introduced via Generalized Uncertainty Principle (GUP) deformed Poisson brackets in the Hamiltonian framework. We show that free-particle geodesics remain unaffected at leading order, but external potentials induce velocity-dependent corrections, implying possible violations of the equivalence principle. As an application, we analyze modified trajectories in Friedmann-Lemaitre-Robertson-Walker (FLRW) universes dominated by dust, radiation, stiff matter, and dark energy. Our results establish a unified approach to conservative, dissipative, and GUP-corrected geodesics, providing a framework to probe the interplay between external forces, spacetime curvature, and Planck-scale physics.
@@
General Relativity (GR) is built on two key ideas. The first is the equivalence principle, which says that gravity acts the same on all objects, no matter their shape or composition. The second is Mach’s principle, which states that matter shapes the geometry of spacetime, and this curvature then guides matter to move along the straightest possible paths, called geodesics and they are crucial for understanding phenomena such as cosmic horizons, dark energy, and gravitational lensing, as well as for studying observational effects like cosmic bubble collisions and theoretical frameworks such as the AdS/CFT correspondence [5, 6, 7, 8, 9, 10, 11].
@@ III Application
To investigate the qualitative behaviour of the modified geodesic equation, we consider the FLRW spacetime in the presence of both conservative external forces and dissipative effects, as discussed in the following subsections.
@@ III.1.2 Example for dissipative Lagrangian
We turn to the question of how dissipative effects impact the geodesic motion of a test particle in an FLRW spacetime. From Eq. (8), the dissipation exponent in the comoving frame with four-velocity $u^{\kappa}=(1,0,0,0)$ is given by
@@
where $H(t)$ denotes the Hubble parameter.
In cosmological settings, dissipative processes can originate from a variety of microphysical and effective mechanisms, such as bulk viscosity, particle production, scalar-field interactions, and non-equilibrium thermodynamic effects. These scenarios generally lead to nontrivial functional forms of the dissipation function $Q(x^{\alpha})$ [70, 71, 72, 73, 74, 75, 76, 77]. To preserve analytical tractability while isolating the qualitative impact of dissipation on geodesic evolution, we adopt the constant parametrisation, which is one of the form proposed in Ref. [16], $Q(x^{\alpha})=Q_{0}=\text{constant}$. This simplification maintains full compatibility with FLRW symmetry and enables a transparent analytical characterisation of the dissipative modifications to the geodesic equation.
In this case, the geodesic equation simplifies to
@@
where the term $\partial_{0}\Gamma=2H(t)Q(t)$. The physical origin of dissipation in this framework is encoded in the scalar function $Q(x^{\alpha})$. It acts as a phenomenological parameter that characterises the exchange of energy between the particle and an external environment, such as a thermal bath, background radiation, or effective degrees of freedom that have been coarse-grained out. The coupling with the spacetime expansion, expressed through $\nabla_{\kappa}u^{\kappa}$, ensures that the dissipative effect depends on the local geometry and flow. The geodesic path will be affected by the values of the dissipation function $Q$, which is chosen to be constant here. For positive value ($Q>0$), it means the energy is dissipating in the surrounding, while negative values of $Q$ correspond to effective energy injection as in dissipative quintessence and inflationary scenarios [16]. The qualitative behaviour of this set of differential equations, Eq.(31,32), through geodesic curve is represented in Fig.2. We show some plot of geodesic curves for different values of $Q$.
@@ IV The Geodesic equations under minimal length effect (GUP)
In the absence of a complete and consistent theory of quantum gravity, a variety of distinct approaches have been developed to capture possible Planck-scale effects. Among them, the Generalized Uncertainty Principle (GUP) provides a phenomenological framework that encodes the existence of a minimal measurable length, thereby modifying the standard Heisenberg uncertainty principle. It can be expressed [44, 46, 69] as
@@
where $x$ and $p$ denote the position and its canonically conjugate momentum, respectively. The deformation parameter $\beta$ encodes the strength of quantum-gravity that becomes significant where the quantum gravitational effects are visible. The parameter $\beta$ typically arises in various approaches to quantum gravity, such as string theory, loop quantum gravity, and doubly special relativity. The dimension of $\beta$ is inverse of the square of Planck momentum as represented as $\beta=\beta_{0}/(m_{P}c)^{2}$, where $m_{P}$ is the Planck mass and $\beta_{0}$ is a dimensionless parameter. The value of $\beta_{0}$ is model-dependent and can be fixed or constrained through phenomenological studies in both quantum-mechanical and classical regimes.
@@
It is crucial to emphasize that Eq. (36) does not represent an arbitrary noncanonical choice of symplectic structure. In conventional Hamiltonian formulations, noncanonical symplectic forms may arise for example, in the quantization of dissipative systems [78] yet such deformations are purely kinematical and can always be removed by a suitable canonical transformation, leaving the phase-space geometry globally flat.
In contrast, the deformation encoded in Eq. (36) originates from minimal measurable length motivated by quantum gravity phenomenology. Here, the symplectic structure depends intrinsically on the momentum coordinates, implying a nontrivial curvature of momentum space [^1] Darboux’s theorem guarantees that any symplectic manifold is locally equivalent to the canonical form; that is, locally one can always find coordinates $(q_{i},p_{i})$ such that $\omega=\sum_{i}dp_{i}\wedge dq_{i}$. This implies that classical noncanonical structures merely reflect coordinate choices rather than intrinsic curvature of phase space. The wedge product provides a coordinate-independent way to define the symplectic two-form, and for GUP-deformed systems, one must check explicitly whether $d\omega=0$ holds [79, 80, 81].. This interpretation establishes a natural link between the GUP framework and the theories of Doubly Special Relativity and relative locality, where Lorentz transformations act nonlinearly on momenta. Therefore, while the GUP algebra may resemble noncanonical formulations used in certain classical systems, their physical content is fundamentally distinct. The GUP deformation reflects an intrinsic modification of phase-space geometry, whereas classical noncanonical structures merely encode coordinate choices within a flat phase space.
It should also be noted that this procedure of taking the GUP-deformed commutator and promoting it directly to a Poisson bracket has limitations. In particular, as discussed in [69], when taking the formal limit $\hbar\rightarrow 0$, the deformation term often diverges as $(1/\hbar)$. Strictly speaking, this signals an inconsistency in deriving the classical limit directly from the quantum commutator. On the other hand, it has been argued in [83] that although the factor $\hbar^{-1}$ diverges in the classical limit, the GUP parameter $\beta$ encodes the strength of quantum-gravity effects, where the dimensionless parameter $\beta_{0}$ can be chosen to be vanishingly small. This ensures that the overall factor $\beta_{0}\hbar^{-1}$ remains small but finite, thereby providing a consistent classical limit.
@@ IV.2 GUP-Corrected Geodesic Equation in the Presence of a Potential
We further extended our analysis by incorporating minimal-length effects through a GUP-deformed symplectic structure at the classical level. Our results show that free-particle geodesics remain unaffected to leading order in the deformation parameter $\beta$, thereby preserving the equivalence principle in the absence of external interactions. However, when external potentials are present, the geodesic equation acquires velocity-dependent corrections proportional to $\beta$. These velocity-dependent modifications in the GUP-modified geodesic equation signal violations of the equivalence principle.
A noncanonical symplectic structure alone is insufficient to uniquely attribute the behavior to quantum-gravity effects. Noncanonical phase-space structures arise naturally in purely classical systems for example, in the constrained Hamiltonian dynamics [78] or in the Hamiltonian description of charged particles in electromagnetic fields [82], in such cases they simply reflect nontrivial coordinate choices on an otherwise flat phase space. By contrast, the GUP-deformed symplectic structure employed in this work is motivated by quantum-gravity phenomenology as it encodes a minimal measurable length and leads to a momentum-dependent deformation of the underlying phase space. Thus, while classical and GUP-induced noncanonical structures may appear algebraically similar, their physical interpretations are fundamentally distinct.
Altogether these findings demonstrate that while pure geodesic motion is robust against Planck-scale modifications, the interplay between external forces, dissipation, and GUP corrections leads to a rich phenomenology, particularly in cosmological backgrounds. Our qualitative analysis of trajectories in dust, radiation, stiff matter, and dark-energy dominated universes illustrates how such effects can alter the geodesic paths at both the classical and quantum-gravity levels, as depicted in Figs. (1, 2, 3).
While we adopt a constant dissipation coefficient $Q_{0}$ to obtain a qualitative analysis, the present framework can be systematically extended to more general forms of $Q(x^{\alpha})$. Adopting dissipation functions parametrized by the Hubble rate, the matter density, or the redshift would facilitate an examination of how distinct dissipative mechanisms affect geodesic trajectories and geodesic deviation in FLRW spacetimes, contributing to a broader phenomenological understanding of dissipative gravitational dynamics.
@@
Appendix A: Derivation of the Generalized Uncertainty Principle
In this Appendix we derive the commonly used form of the Generalized Uncertainty Principle (GUP) starting from the modified canonical commutator
$x,p]=i\hbar\,[\,1+\beta p^{2}\,],$ (54)
which appears in several phenomenological models of Planck-scale physics [44, 46, 69]. The Robertson (or Robertson-Schrödinger) uncertainty relation for two self-adjoint operators $A$ and $B$ states
$\Delta A\,\Delta B\;\geq\;\frac{1}{2}\bigl|\langle[A,B]\rangle\bigr|,$ (55)
considering the operators as $A=x$ and $B=p$ and using Eqns.(54,55) gives
$\Delta x\,\Delta p\;\geq\;\frac{1}{2}\bigl|\langle[x,p]\rangle\bigr|=\frac{\hbar}{2}\bigl\langle 1+\beta p^{2}\bigr\rangle=\frac{\hbar}{2}\bigl(1+\beta\langle p^{2}\rangle\bigr).$ (56)
Now express $\langle p^{2}\rangle$ in terms of the momentum variance and mean,
$\langle p^{2}\rangle=(\Delta p)^{2}+\langle p\rangle^{2}\;\geq\;(\Delta p)^{2},$ (57)
so Eq.(56) implies the commonly quoted GUP bound
$\Delta x\,\Delta p\;\geq\;\frac{\hbar}{2}\bigl(1+\beta(\Delta p)^{2}\bigr)\;.$ (58)
Equation (58) makes explicit that large $\Delta p$ increases the lower bound on $\Delta x$, leading to a nonzero minimum position uncertainty. To find the minimal $\Delta x$, rewrite the Eq.(58) as
$\Delta x\;\geq\;\frac{\hbar}{2}\Bigl(\frac{1}{\Delta p}+\beta\,\Delta p\Bigr),$
and minimize the right-hand side with respect to the positive variable $\Delta p$. The derivative vanishes at
$\frac{1}{(\Delta p)^{2}}+\beta=0\quad\Rightarrow\quad\Delta p_{\rm crit}=\frac{1}{\sqrt{\beta}}.$
Substituting back gives the minimal uncertainty in position as
$\Delta x_{\rm min}=\hbar\sqrt{\beta}.$ (59)
Thus, the modified commutation relation (54) implies the existence of a minimal measurable length $\Delta x_{\rm min}=\hbar\sqrt{\beta}$.
@@ Acknowledgements.
Part of this work was carried out during G.B.’s visit to IUCAA, Pune, and G.B. thanks Prof. Ranjeev Mishra for support during the stay. SDP acknowledges the indirect support of the global community of researchers, institutions, and organizations dedicated to advancing fundamental physics, which makes work of this kind possible. The work of M.K. was performed in Southern Federal University with financial support of grant of Russian Science Foundation № 25-07-IF.